# Data Augmentation is Dead, Long Live Data Augmentation

### v1

@@ Abstract
Textual data augmentation (DA) is a prolific field of study where novel techniques to create artificial data are regularly proposed, and that has demonstrated great efficiency on small data settings, at least for text classification tasks. In this paper, we challenge those results, showing that classical data augmentation is simply a way of performing better fine-tuning, and that spending more time fine-tuning before applying data augmentation negates its effect. This is a significant contribution as it answers several questions that were left open in recent years, namely : which DA technique performs best (all of them as long as they generate data close enough to the training set as to not impair training) and why did DA show positive results (facilitates training of network). We furthermore show that zero and few-shot data generation via conversational agents such as ChatGPT or LLama2 can increase performances, concluding that this form of data augmentation does still work, even if classical methods do not.
Keywords: NLP, Classification, Data Augmentation, ChatGPT, Llama2
[…]
Data Augmentation is Dead, Long Live Data Augmentation
Frédéric Piedboeuf, Philippe Langlais
Université de Montréal, RALI
frederic.piedboeuf@umontreal.ca
Abstract content
1.   Introduction
@@
Data augmentation (DA) consists in generating artificial data points with the hope of improving the training of a model. More specifically, interpretable textual DA takes interest into whether new sentences can increase performances, with the most common task for testing algorithms being text classification. Research on the subject generally finds that DA provides a great increase on small data settings Chen et al. (2021); Kumar et al. (2021), a small increase in classification with medium datasets (up to 1000 examples) Karimi et al. (2021); Liesting et al. (2021), and almost no increase on large datasets Kobayashi (2018); Yang et al. (2020).
DA algorithms proposed in the literature have an important effect not only because they guide practitioners towards one technique or the other, but also because they contribute to the understanding of how neural networks learn. Showing, for example, that a word substitution algorithm helps a pretrained model like BERT has important implications about its inner working.
@@
In this paper, we show that inadequate fine-tuning of the baselines (model trained without DA) from previous studies lead to an overestimation of the impact of textual DA. In fact, we show that by simply playing with the patience parameter as well as occasionally adding label smoothing Szegedy et al. (2015), we can raise the performance of the baseline classifier to be as good as performances with DA, and that when the classifier is trained correctly DA does not bring improvements. We further compare novel DA methods using ChatGPT and Llama2 and show that the only thing that can consistently increase classification performance is generating data akin to external data (with zero or few-shot data generation), rather than generating new data from the current training distribution (by paraphrasing or modifying a sentence from the training set), as classical DA approaches do. Finally, we consider more realistic approaches to DA, where we either don’t have access to clean validation data, or where we adjust the training/validation split to better reflect real-world conditions.
@@
Overall, this paper covers several important contributions. The first and most important of them lies in showing that classical data augmentation (DA methods from before the advent of Very Large Language Models, or VLLMs) does not work on textual classification, and we also answer why previous studies showed positive results. These positive results have been a big question in recent years, as nobody could explain with satisfaction why generated sentences helped transformers learn better, the best hypothesis being that it brought some kind of regularization to the network (Feng et al., 2021; Queiroz Abonizio and Barbon Junior, 2020).
The second contribution is showing that by using VLLMs to generate novel sentences (and not just paraphrase or modify existing ones), the performance increases, but that this is still not as efficient as collecting and annotating novel data. We furthermore analyze two strategies for generating novel data, namely zero-shot data generation (using the task description) and 3-shot data generation (giving three examples to the VLLM), showing that in most case the 3-shot strategy largely outperforms zero-shot generation.
The third contribution joins a small but growing literature comparing VLLMs in showing that ChatGPT outperforms Llama2 (Liu et al., 2023; Plátek et al., 2023; Guo et al., 2023). To our knowledge, we are the first to study this in the context of data augmentation.
@@
Finally, our last contribution is a questioning of the use of the validation data in data augmentation for small data settings. DA research often supposes the availability of clean validation samples, but its usage is not realistic in most cases. With the splits used in those studies, one finds itself with a few data points for training the model (generally less than 30), and a few hundreds for validation. In this paper, we attempt to consider better splitting and more realistic use of data.
@@
The paper is organized as such. In Section 2, we go over the literature on interpretable textual DA. Sections 3 and 4 explain the strategies, protocol, and datasets used in this paper. Then Section 5 presents the results and Section 6 go over analysis, including the evaluation of more realistic uses of the validation data. We finally include a discussion about the meaning of our work in Section 7.
2.   Related Work
@@
We may categorize DA techniques in three broad families : word-level operations, paraphrasing, and generative DA. Furthermore, we refer to DA techniques that were developed before VLLMs as classical DA.
@@
The first family creates variations of existing sentences by affecting individual words. The most seminal technique of this category is EDA, which considers four operations : word substitution, word deletion, word swapping, and insertion of related vocabulary (Wei and Zou, 2019; Liesting et al., 2021). For insertion and substitution, EDA uses WordNet to find synonyms of words from the input sentences. Another simple but efficient word-level technique is AEDA Karimi et al. (2021), where random punctuation is inserted in between words of the sentence.
@@
Other ways of replacing words have also been considered, such as the use of a pre-trained neural network to predict masked words Kobayashi (2018), using pre-trained embeddings to find words close in the embedded space Marivate and Sefara (2020), or using BERT/BART conditioned on the class to predict masked words or spans Wu et al. (2019); Kumar et al. (2021). These techniques are generally easy to implement but bring little diversity to the dataset.
@@
The second family of techniques acts at the sentence level, by considering the whole sentence for creating paraphrases. The seminal technique representing this family of techniques is Back-translation (BT), in which a sentence is translated to another language and then back into English (Hayashi et al., 2018; Yu et al., 2018; Corbeil and Ghadivel, 2020; Edunov et al., 2018). The use of generative models for paraphrasing has also been considered, for example by encoding and decoding a sentence through a VAE (Mesbah et al., 2019; Yerukola et al., 2021; Nishizaki, 2017). Some other strategies that have been considered are modifying syntactic trees Coulombe (2018), fine-tuning BART on an in-domain corpus of paraphrases Okur et al. (2022), using an off the shelf T5 model for paraphrasing (T5-Tapaco) Piedboeuf and Langlais (2023), or using VLLMs such as ChatGPT for paraphrasing Fang et al. (2023).
@@
Finally, generative methods aim to generate novel sentences from the same distribution as the training data. With an ideal generative strategy, this is equivalent to collecting new data, without the annotation cost associated to the collection process. Early studies of these techniques used GPT-2 (Kumar et al., 2020; Liu et al., 2020; Yang et al., 2020; Bayer et al., 2023) or VAEs Qiu et al. (2020); Piedboeuf and Langlais (2022), but recently focus has shifted to the use of VLLMs, first with GPT3 (Yoo et al., 2021; Sahu et al., 2022), and then with ChatGPT for zero or small data generation (Møller et al., 2023; Ubani et al., 2023; Shushkevich and Cardiff, 2023; Sharma and Feldman, 2023).
@@
Some recent works pertinent to this paper are the research of Kim et al. (2022); Zhu et al. (2023)Kim et al. (2022) question the training-validation splitting methodology in semi-supervised learning, showing that it is more efficient to fine-tune on augmented samples (created with DA) and use the original training sentences as validation data, instead of using a classical training/validation split. In Zhu et al. (2023), the authors note that weakly supervised learning approaches are evaluated by assuming the availability of clean validation samples, which is not often the case when working with small data. They notably develop novel methods for testing training under more realistic small data learning settings. In Section 6, we take inspiration of those papers to define our new strategies.
3.   Data augmentation strategies
Kumar et al. (2021); Piedboeuf and Langlais (2023) compared several DA methods, from which we select strategies that have been shown to be efficient. Concretely, we test EDA, AEDA, CBERT, CBART, T5, BT, and three strategies with VLLMs : paraphrasing, zero-shot generation, and k-shot generation. All strategies are illustrated in Figure 1.
EDA (multiple operations on the words) and AEDA (insertion of punctuation) can be implemented fairly simply. Following experiments and results from the literature, we affect 10% of the words of a sentence in EDA, and use the formula given in Karimi et al. (2021) to calculate the number of punctuation signs to insert for AEDA.
CBERT and CBART both use the task of masked words prediction to generate new words conditionally on the class. Concretely, we mask words, prepend the class to the sentence (with a separation token), and fine-tune the model to predict these masked words. For generating new sentences, we follow the same process. The difference between CBERT and CBART is that CBART can predict spans of words instead of just individual tokens, allowing it more flexibility in the generated sentences.
Both BT and T5-Tapaco aim to produce paraphrases of the original sentences to bring diversity to the training set. In BT, paraphrases are generated by translating the sentence into a second language and then back into English, and we use WMT [^1] https://huggingface.co/facebook/wmt19-De-en with German as a pivot language, which has shown good performances in the past Edunov et al. (2018). T5-Tapaco makes use of T5-small-Tapaco [^2] https://huggingface.co/hetpandya/t5-small-tapaco, which is a T5 model fine-tuned on the TaPaCo paraphrase corpus (Scherrer, 2020), allowing us to directly generate paraphrases of the sentences.
In terms of VLLMs, we test three strategies, which are the three standards strategies we could find in the literature : paraphrasing existing sentences, zero-shot generation by giving the VLLMs a description of the task, and a k-shot generation strategy where we give examples of the given class. We give the exact prompts used in appendix A. Furthermore, we test the use of Llama2 (using LLama-2-13B-Chat) with the same prompts. We denote in this paper these strategies by the letters P (paraphrase), ZS (zero-shot), and 3S (3-shot generation). We opted on 3-shot for the K-shot generation strategy based on results from exploratory experiments.
Figure 1: Strategies tested in this paper.
Finally, we implement three simple baselines. The first one consists in training the network without the use of DA (denoted “Baseline”), and the second one is an idealized strategy (denoted “‘Perfect”) where we fetch additional unused data from the training set to act as generated data. This gives an idea of the results obtainable should one collect and annotate data instead of generating new data.
The last baseline is a strategy which we denote “Copy”, and which artificially inflates the size of the training set by copying multiple times the original data, without modification. If the only effect of DA is truly to help fine-tune the network better, and the modification brought by the DA strategies are not helpful for that, then the copy strategy should be just as efficient as all other classical data augmentation strategies. We give the code with all hyperparameters in the associated github. [^3] Github will be made available upon acceptance of the paper.
4.   Datasets and protocol
@@
We use five datasets to test the various data augmentation strategies : SST-2, Irony, IronyB, TREC6, and SNIPS. SST-2 (Socher et al., 2013) is a dataset of movie review classification with the binary classes of positive or negative. Irony and IronyB (Van Hee et al., 2018) are the binary and multiclass version of an Irony detection dataset. In the binary task (Irony), one must detect if tweets are ironic or not, and in the multiclass version one must as well determine which type of Irony the tweet represents, if the tweet is ironic (between polarity clash, situational irony, and other irony). TREC6 (Li and Roth, 2002) is a task of question classification, where questions must be separated into six classes (abbreviation, description, entities, human beings, locations, and numeric values). Finally, SNIPS Coucke et al. (2018) is a dataset of intent classification, where short commands have to be classified into different intent (PlayMusic, GetWeather, RateBook, SearchScreeningEvent, SearchCreativeWork, AddToPlaylist, BookRestaurant). Some characteristics of the datasets are available in Table 1.
Table 1: Characteristics of the tasks tackled in this study. The length of the sentences is defined by the number of white-space separated tokens.
@@
To define our training protocol, we first look at the literature to understand what is usually done. Piedboeuf and Langlais (2023) use variable patience for fine-tuning, depending on the dataset and dataset size, between 5 and 20 epochs, and Kumar et al. (2021) use 100 epochs of warmup followed by 8 epochs of fine-tuning from which they select the best model. Wei and Zou (2019) use early stopping with patience of 3 epochs on CNN and RNNs.
In contrast, we use a patience parameter of 50, using the validation set to find when to stop, and we fine-tune the learning rate and the label smoothing (a mean of introducing noise to the labels for regularization) before applying data augmentation, using grid search. We define here the patience as the number of epochs during which the validation performance doesn’t increase before we stop the training process, and we then select the best model from previous epochs. We report accuracy for binary tasks and the macro-F1 for multiclass tasks.
Our focus in this paper is data augmentation for small data learning, for which we use starting size of 10 and 20, but we also report results for medium training sizes (500 and 1000) in Appendix B, a setting which is commonly explored in the literature. When subsampling to create the training set, we make sure to choose an equal number of data points for each class, by going over the fixed dataset size if needed (for example, the actual dataset size for TREC6 and starting size of 10 is 12; two sentences per class).
@@
The final parameter is the ratio of generated-to-genuine examples. There seems to be no strict rules about this ratio, but we found in both the literature and experiments that smaller sizes of datasets benefited more from larger ratios, while larger starting sizes did not show improvements with ratios larger than one. Therefore, we use a ratio of 10 for the small data settings and of one for the larger dataset sizes. [^4] Based on our main results, we can hypothesize that this is simply due to pretrained models needing more time to be fine-tuned correctly on smaller datasets than medium or large ones, and the larger ratio leads to a larger number of batches, meaning more training time.
Something worth noting is that all of these protocols (including ours) use the validation data to fine-tune both the DA algorithm and the model to train. This is somewhat questionable, as keeping only 10 or 20 data points out of a few hundreds might not be the best way to utilize the data. While we keep this protocol in the main section to stay in accordance with the DA literature, we propose more realistic evaluation protocol of DA in Section 6.
The norm in data augmentation literature is to give the average and standard deviation over a fixed number of runs. However, to confirm whether the occasional improvements are due to randomness of the model or the effect of DA, we include in this paper statistical tests.
Table 3: Average metric over 15 runs for the training set sizes of 10 (left) and 20 (right) with a ratio of 1. We report accuracy for binary tasks and macro-f1 for multiclass ones. STDs are between 0.6 and 3.0, depending on the dataset. Stars represent results for which the difference with the baseline was found to be statistically significant.
SST2 Irony IronyB Trec6 SNIPS Average
Perfect 82.3*/85.8* 62.4*/58.7 19.3/33.7 72.5*/80.6* 94.0*/94.9* 66.1/70.7
Copy 61.0/68.1 54.4/57.1 25.5/30.3 36.4/49.3* 77.7/83.4* 51.0/57.6
EDA 63.4/69.0 54.9/58.9 23.8/30.1 35.8/48.0* 76.6/83.2* 50.9/57.8
AEDA 62.4/67.1 55.4/56.9 23.4/27.4 33.2/45.7 78.0/84.0* 50.5/56.2
BT 60.2/66.8 55.1/58.0 24.7/27.9 29.1/51.5* 75.0/82.6* 48.8/57.4
CBERT 62.5/65.4 56.3*/58.5 23.2/28.3 36.4/44.5 76.3/83.4* 50.9/56.0
CBART 62.5/65.6 54.5/57.7 23.6/29.0 36.9/47.2 77.0/84.5* 50.9/56.8
T5 64.1/67.4 56.3/58.3 24.1/28.1 34.3/48.5* 76.7/84.1* 51.1/57.2
GPT3.5-P 65.0*/68.3 56.6*/57.5 22.7/28.3 34.0/40.5 80.2/84.7* 51.7/55.9
GPT3.5-ZS 82.0*/78.7* 56.4*/56.1 23.5/24.6 40.0*/48.1* 84.8*/88.7* 57.3/59.2
GPT3.5-3S 87.7*/76.1* 49.2/53.6 20.9/24.6 58.4*/63.9* 87.2*/89.6* 60.7/61.5
Llama2-P 66.3*/68.6 55.3/57.1 24.6/24.6 33.4/41.6 78.1/82.8* 51.6/55.0
Llama2-ZS 74.8*/74.0* 54.3/57.2 22.2/24.1 37.3/47.3 78.1/84.1* 53.4/57.3
Llama2-3S 63.3/64.4 54.1/54.8 21.0/27.2 48.6*/55.8* 83.3*/88.2* 54.0/58.1
For each starting size (10, 20) and dataset, we perform a two-tailed paired t-test, comparing the algorithms to each other [^5] We choose to do this instead of a Hotelling’s T-square to take into account the fact the DA might work on some datasets but not others. For example, one would intuitively expect synonym replacement to work better on sentiment classification than on Irony detection.. As is standard, we assess the two distributions to be different if the p-value is lower than 0.05.
5.   Results

### v2

# On Evaluation Protocols for Data Augmentation in a Limited Data Scenario

@@ Abstract
Textual data augmentation (DA) is a prolific field of study where novel techniques to create artificial data are regularly proposed, and that has demonstrated great efficiency on small data settings, at least for text classification tasks. In this paper, we challenge those results, showing that classical data augmentation (which modify sentences) is simply a way of performing better fine-tuning, and that spending more time doing so before applying data augmentation negates its effect. This is a significant contribution as it answers several questions that were left open in recent years, namely : which DA technique performs best (all of them as long as they generate data close enough to the training set, as to not impair training) and why did DA show positive results (facilitates training of network). We further show that zero- and few-shot DA via conversational agents such as ChatGPT or LLama2 can increase performances, confirming that this form of data augmentation is preferable to classical methods.
On Evaluation Protocols for Data Augmentation
in a Limited Data Scenario
Frédéric Piedboeuf Université de Montréal, RALI frederic.piedboeuf@umontreal.ca                        Philippe Langlais Université de Montréal, RALI
1 Introduction
@@
Data augmentation (DA) consists in generating artificial data points with the hope of improving the training of a model. In this paper, we focus on interpretable textual DA (methods that generate new sentences) for text classification in a limited data scenario, a practical setting of interest Chen et al. (2021). Research generally finds that DA provides a great increase on small data settings Chen et al. (2021); Kumar et al. (2021), a small increase in classification with medium datasets (up to 1000 examples) Karimi et al. (2021); Liesting et al. (2021), and almost no increase on large datasets Kobayashi (2018); Yang et al. (2020); Okimura et al. (2022).
@@
In this paper, we show that existing experimental protocols in DA studies are misleading. In particular, inadequate fine-tuning of the baseline (model trained without DA) from previous studies lead to an overestimation of the impact of textual DA, and training the models for longer results in an absence of gains with previous DA methods. We also consider more realistic protocols for DA evaluation, where we either don’t have access to clean validation data, or where we adjust the training/validation split to better reflect real-world conditions. Lastly, we compare newer DA methods using ChatGPT Ouyang et al. (2022) and Llama2 Touvron et al. (2023) and show that the only thing that can consistently increase classification performance is generating data akin to external data (with zero or few-shot data generation), rather than generating new data from the current training distribution (by paraphrasing or modifying training sentences), as classical DA approaches do.
@@
Overall, this paper covers several important contributions. The first and most important of them lies in showing that classical data augmentation (DA methods from before the advent of very Large Language Models, or LLMs) does not work on textual classification, and we also answer why previous studies showed positive results. These positive results have been a big question in recent years, as nobody could explain with satisfaction why generated sentences helped transformers learn better, the best hypothesis being that it brought some kind of regularization to the network (Feng et al., 2021; Queiroz Abonizio and Barbon Junior, 2020).
@@
Our second contribution is a questioning of the use of the validation data in data augmentation for small data settings. DA research generally supposes the availability of clean validation samples, an often unrealistic setting. With the splits used in those studies, one finds themself with a few tens of data points for training the model and a few hundreds for validation. In this paper, we consider better splitting and more realistic uses of data, showing that it is often more advantageous to use all available data as training data and to fine-tune for a longer time.
The third contribution is showing that by using LLMs to generate novel sentences (and not just paraphrasing existing ones), the performance increases, but that this is still not as efficient as manually collecting and annotating novel data. We furthermore analyze two strategies for generating novel data, namely zero-shot data generation (using the task description) and 3-shot data generation (giving three examples to the LLM), showing that in most cases the 3-shot strategy largely outperforms zero-shot generation.
The last contribution joins a small but growing literature comparing LLMs in showing that ChatGPT outperforms Llama2 (Liu et al., 2023; Plátek et al., 2023; Guo et al., 2023). To our knowledge, we are the first to study this in the context of DA.
@@
The paper is organized as such. Section 2 goes over the literature on interpretable textual DA. Sections 3 and 4 explain respectively the datasets and the DA methods we compare. Section 5 describes our evaluation protocols, then Section 6 presents our results. We discuss our work in Section 7.
2 Related Work
@@
We may categorize DA techniques in three broad families : word-level operations, paraphrasing, and generative DA. We note that this is not an absolute categorization, and each of these families regroup several families of techniques. We also refer to techniques that were developed before LLMs as classical DA. These modify the starting sentences in some ways, bringing variations to the dataset.
@@
The first family does so by affecting individual words. The most seminal technique of this category is EDA, which considers four operations : word substitution, word deletion, word swapping, and insertion of related vocabulary (Wei and Zou, 2019; Liesting et al., 2021). For insertion and substitution, EDA uses WordNet Miller (1998) to find synonyms of words from the input sentence. Another simple word-level technique is AEDA Karimi et al. (2021), where random punctuation is inserted in between words of the sentence.
@@
Other ways of replacing words have also been considered, such as the use of a pre-trained neural network to predict masked words Kobayashi (2018), using pre-trained embeddings to find words close in the embedded space Marivate and Sefara (2020), or using BERT/BART conditioned on the class to predict masked words or spans Wu et al. (2019); Kumar et al. (2021). While easy to implement, those techniques tend to bring little diversity.
@@
The second family of techniques acts at the sentence level, by considering the whole sentence for creating paraphrases. The seminal technique representing this family is Back-translation (BT), in which a sentence is translated to another language and then back into English (Hayashi et al., 2018; Yu et al., 2018; Corbeil and Ghadivel, 2020; Edunov et al., 2018). The use of generative models for paraphrasing has also been considered, for example by encoding and decoding a sentence through a VAE (Mesbah et al., 2019; Yerukola et al., 2021; Nishizaki, 2017). Some other strategies that have been considered are modifying syntactic trees Coulombe (2018), fine-tuning BART on an in-domain corpus of paraphrases Okur et al. (2022), using an off the shelf T5 model for paraphrasing (T5-Tapaco) Piedboeuf and Langlais (2023), or using LLMs such as ChatGPT for paraphrasing Fang et al. (2023).
@@
Finally, generative methods aim to generate novel sentences from the same distribution as the training data. With an ideal generative strategy, this is equivalent to collecting new data, without the annotation cost associated to the collection process. Early studies of these techniques used GPT-2 (Kumar et al., 2020; Liu et al., 2020; Yang et al., 2020; Bayer et al., 2023) or VAEs Qiu et al. (2020); Piedboeuf and Langlais (2022), but recently focus has shifted to the use of LLMs, first with GPT3 (Yoo et al., 2021; Sahu et al., 2022), and then with ChatGPT for zero or small data generation (Møller et al., 2023; Ubani et al., 2023; Shushkevich and Cardiff, 2023; Sharma and Feldman, 2023).
@@
Some recent works pertinent to this paper are the research of Kim et al. (2022); Zhu et al. (2023)Kim et al. (2022) question the training-validation splitting methodology in semi-supervised learning, showing that it is more efficient to fine-tune on augmented samples (created with DA) and use the original training sentences as validation data, instead of using a classical training/validation split. In Zhu et al. (2023), the authors note that weakly supervised learning approaches are evaluated by assuming the availability of clean validation samples, which is not often the case when working with small data. They notably develop novel methods for testing training under more realistic small data learning settings. In Section 5, we take inspiration of those papers to define our new protocols.
3 Datasets
@@
We use five popular datasets to test the various data augmentation strategies : SST-2, Irony, IronyB, TREC6, and SNIPS. SST-2 (Socher et al., 2013) is a dataset of movie review classification with the binary classes of positive or negative. Irony and IronyB (Van Hee et al., 2018) are the binary and multiclass version of an Irony detection dataset. In the binary task (Irony), one must detect if tweets are ironic or not, and in the multiclass version one must as well determine which type of Irony the tweet represents, if the tweet is ironic (between polarity clash, situational irony, and other irony). TREC6 (Li and Roth, 2002) is a task of question classification, where questions must be separated into six classes (abbreviation, description, entities, human beings, locations, and numeric values). Finally, SNIPS Coucke et al. (2018) is a dataset of intent classification, where short commands have to be classified into different intents such as PlayMusic or GetWeather. Some characteristics of the datasets are available in Table 1.
@@
Table 1: Characteristics of the classification tasks tackled in this study. The length of the sentences is defined by the number of white-space separated tokens.
Due to lack of computing budget [^1] As we carefully fine-tune every baseline and run every experiment 15 times, this demands a significant time investment. we leave the investigations of other tasks such as those considered in Chen et al. (2021) to future work.
4 Data augmentation methods
While DA methods are continuously being proposed, objective evaluation is difficult due to the lack of extensive comparison studies. We point the reader to Chen et al. (2021) for a literature review on DA methods, and Ding et al. (2024) for one specific to generative approaches. Here, we compare LLM based methods to previous (classical) approaches. All strategies are illustrated in Figure 1. Code and all hyperparameters are available in the additional material.
4.1 Classical methods
Kumar et al. (2021); Piedboeuf and Langlais (2023) compared several “classical” DA methods, from which we select three families of strategies that have been shown to be efficient. Concretely, we test word-manipulation methods (EDA, AEDA), conditional contextual based methods (CBERT, CBART), and paraphrase based methods (T5, BT).
EDA (multiple operations on words) and AEDA (insertion of punctuation) can be implemented simply and are resource efficient. Following experiments and results from the literature, we affect 10% of the words of a sentence in EDA, and use the formula given in (Karimi et al., 2021) to calculate the number of punctuation signs to insert for AEDA. [^2] EDA works by randomly selecting one of four operations (insertion of related words, word swapping, word deletion, and word substitution) and applying it to a percentage of the words of the sentence. AEDA works by simply inserting random punctuation (among ”?”, ”.”, ”;”, ”:”, ”!”, and ”,”) into the sentence.
CBERT and CBART are more involved methods that leverage the masked words prediction task to generate new words conditionally on the class. Concretely, we mask words, prepend the class to the sentence (with a separation token), and fine-tune the model to predict the masked words. Generation of sentences follow the same process. The difference between CBERT and CBART is that latter can predict spans of words, allowing it more flexibility in the generated sentences.
BT and T5-Tapaco aim to produce paraphrases of the original sentences to bring diversity to the training set. In BT, paraphrases are generated by translating the sentence into a second language and then back into English, and we use WMT [^3] https://huggingface.co/facebook/wmt19-De-en with German as a pivot language, which has shown good performances in the past Edunov et al. (2018). T5-Tapaco makes use of T5-small-Tapaco [^4] https://huggingface.co/hetpandya/t5-small-tapaco, which is a T5 model fine-tuned on the TaPaCo paraphrase corpus (Scherrer, 2020), allowing us to directly generate paraphrases of the sentences.
4.2 Large Language Models
LLMs are ubiquitous in NLP, and have rapidly gained traction in data augmentation. Here, we test three standard strategies : paraphrasing sentences (P), zero-shot generation (ZS) by giving the LLMs a description of the task, and a 3-shot (3S) generation strategy where we give examples of the given class in addition to the description. Exact prompts used are provided in appendix B. We test two models, ChatGPT and Llama2 (using LLama-2-13B-Chat) with the same prompts.
Note that we only use LLMs to generate new data with which we train a classifier, as with any other method we compare. Using LLMs directly to label examples may lead to better classification for some tasks, but we leave this for future investigations. In any case, training a classifier is a practical solution for many problems of interest, as well as a cheaper option to deploy.
Figure 1: Strategies tested in this paper. Green algorithms are Contextual-based methods, yellow are paraphrasing methods, red are word-manipulation methods, and purple are methods using LLMs.
4.3 Baselines
Finally, we implement three simple baselines. The first one consists in training the network without the use of DA (denoted “Baseline”). The second one is an idealized strategy (denoted “‘Perfect”) where we fetch additional unused data from the training set to act as generated data. This gives an idea of the results obtainable should one collect and annotate data instead of using DA.
The last baseline is a strategy which we denote “Copy”, and which artificially inflates the size of the training set by copying multiple times (according to the ratio parameter) the original data, without modification. If the only effect of DA is — as we suspect — to help fine-tune the network better, and the modifications brought by DA are not helpful for that, then the Copy strategy should be just as efficient as all other classical DA methods.
5 Experimental setups
@@
In this section we describe the experimental setups for our two main experiments, that is to say the testing of DA methods with better fine-tuning, and the development of better experimental protocols for DA. Our focus in this paper is data augmentation for small data, for which we use starting size of 10 and 20, but we also report results for medium training sizes (500 and 1000) in Appendix C, a setting commonly explored in the literature. When subsampling to create the training set, we make sure to choose an equal number of data points for each class, by sampling more data if needed (e.g. the actual dataset size for TREC6 and starting size of 10 is 12; two sentences per class). There is no consensus regarding the ratio of generated-to-genuine examples, but we found in both the literature and experiments that smaller datasets benefited more from larger ratios, while larger for larger datasets, we did not observe improvements with ratios larger than one. Thus, we use a ratio of 10 for the small data settings and of one for the larger dataset sizes. [^5] Based on our main results, we can hypothesize that this is simply due to pretrained models needing more time to be fine-tuned correctly on smaller datasets than medium or large ones, and the larger ratio leads to a larger number of batches, meaning more training time.
5.1 Better fine-tuning
To define our training protocol, we look at the literature to understand what is usually done. Piedboeuf and Langlais (2023) use variable patience [^6] The number of epochs during which the validation performance doesn’t increase before we stop the training process, and we then select the best model from previous epochs. for fine-tuning, depending on the dataset and dataset size, between 5 and 20 epochs, and Kumar et al. (2021) use 100 epochs of warmup followed by 8 epochs of fine-tuning from which they select the best model. Wei and Zou (2019) use early stopping with patience of 3 epochs on CNN and RNNs.
In contrast, we use a patience parameter of 50, using the validation set to find when to stop, and we fine-tune the learning rate and use label smoothing Szegedy et al. (2015) (a mean of introducing noise to the labels for regularization) before applying data augmentation, using grid search. [^7] While label smoothing was introduced in this paper as a mean of helping better learning of multiclass tasks, preliminary ablations experiments point that it raises the performance of both augmented and non-augmented dataset equally, and that the conclusions of this paper would be the same with only changing the number of epochs and hyper-parameter tuning. We report accuracy for binary tasks and the macro-F1 for multiclass tasks.
5.2 More realistic uses of data
In academic papers, DA is often evaluated with the assumption that we have access to clean validation data, often in larger quantity than what is available for the training set. This validation data is used to fine-tune both the DA algorithm and the classifier, resulting in unrealistic settings for practitioners, who may not have access to validation data or want to use their data in better ways.
We aim to test more realistic settings with the validation data, inspired by Kim et al. (2022); Zhu et al. (2023), who study the same problems but in different settings (semi-supervised learning and weakly supervised learning) and then discuss the significance of the results for textual DA. While Kim et al. (2022); Zhu et al. (2023) findings are relevant to ours, our interest in the results differ. In both papers, the authors attempt to find the best way to use the available data. In our case, we test several settings we think are more realistic, but or interest is to know if DA useful when there is little or no validation data for fine-tuning the augmentation method and the classifier.
We redefine the train/val. split, using different strategies illustrated in Figure 2. Firstly, we assume the validation data is not available (noVal), and we only have training data. In this setting, we train for a random number of epochs between 50 and 150, which is the range for which our models performed best in small data learning. This is inspired by the protocol of Zhu et al. (2023), where they randomly select a set of hyperparameters when no validation data is available. Secondly, we assume the presence of validation data but redefine the train/val. split, by either 1- using it all as training data and training for a random number of epochs between 4 and 8 (valAsTrain), or 2- keeping some of it as validation data (10, 250) to fine-tune both DA algorithms and the classifier (10 val/250 val).
6 Experiments
In what follows, we add the augmented material to the (small) training set and fine-tune a BERT-Base model [^8] We leave the study of other classifiers for future work. in a supervised way for each classification task, a solution which has been demonstrated efficient Devlin et al. (2019).
6.1 On the need to better fine-tuning